Find the values of the six trigonometric functions of θ with the given constraint. (If an answer is undefined, enter UNDEFINED.)
Function Value       Constraint
sin θ = 0    	     sec θ = −1

sin θ	=	 
cos θ	=	 
tan θ	=	 
csc θ	=	 
sec θ	=	 
cot θ	=

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

Find the values of the six trigonometric functions of θ with the given constraint. (If an answer is undefined, enter UNDEFINED.)
Function Value Constraint
sin = 0 sec = -1
-------------------
cos = 1/sec = -1
================
tan = sin/cos
cot = 1/tan
sec = 1/cos
csc = 1/sin

Answer by solver91311(24713)







You have to choose the sign based on other information given.









Fractions with zero denominators are undefined.

John

My calculator said it, I believe it, that settles it